find three consecutive integers whose sum is 11 more than twice the greatest of the three integers.
Answer by ReadingBoosters(3246)

Set the equation
x + x+1 x+2 = 11 + 2(x+2)
Combine like terms and distribute the 2 on the right
3x + 3 = 11 + 2x + 4
3x - 2x + 3 - 3 = 11 + 2x - 2x + 4 - 3
x = 11 + 4 - 3
x = 12
Integers are 12, 13 and 14
Proof
12 + 13 + 14 = 39
11 + 2(12+2) = 11 + 2(14) = 11 + 28 = 39